a particle is projected with speed 4 2 m/s at an angle 45° with horizontal . after some time t its radius of curvature of path becomes (25/8)m. the value of t

Grade:12

1 Answers

Ashutosh
6 years ago
Radius= u^2 sin 2*Φ/g = (4√2)^2 sin 2*45/g = 32* sin90/g = 32*1/g25/8=32/gImplies that g=8*32/25T= 2u sinΦ/g = 2*4√2 sin45/g = 8√2 *1/√2 *1/g =8*25/(8*32) =25/32~0.7s